OpenStudy (anonymous):

How can you solve linear equations and inequalities in one variable, with equations and co-effecients represented with letters?

OpenStudy (anonymous):

Suppose, ax+3=9 Where x=variable and a=coefficient, Then, Subtracting 3 from both sides, ax+3-3=9-3 or, ax=6 Again dividing both sides by co-efficient a x=6/a Solved!